On December 31, 2020, Dow Steel Corporation had 680,000 shares of common stock and 308,000 shares of 8%, noncumulative, nonconvertible preferred stock issued and outstanding. Dow issued a 4% common stock dividend on May 15 and paid cash dividends of $480,000 and $77,000 to common and preferred shareholders, respectively, on December 15, 2021.

On February 28, 2021, Dow sold 54,000 common shares. In keeping with its long-term share repurchase plan, 4,000 shares were retired on July 1. Dow's net income for the year ended December 31, 2021, was $2,500,000. The income tax rate is 25%.

As part of an incentive compensation plan, Dow granted incentive stock options to division managers at December 31 of the current and each of the previous two years. Each option permits its holder to buy one share of common stock at an exercise price equal to market value at the date of grant and can be exercised one year from that date. Information concerning the number of options granted and common share prices follows:

Options Granted
Date Granted (adjusted for the stock dividend) Share Price
December 31, 2019 15,000 $ 16
December 31, 2020 10,000 $ 25
December 31, 2021 13,500 $ 24


The market price of the common stock averaged $24 per share during 2021.

Required:
Compute Dow's earnings per share for the year ended December 31, 2021. (Do not round intermediate calculations. Enter your answers in thousands. Round "Earnings per share" answers to 2 decimal places.)
Answer #1

Computation of earnings per share is:

Earnings per share = (Net income - Preferred dividend) / Number of weighted shares

= ($2,500,000 - $77,000) / 752,000

= $2,423,000 / 752,000

= 3.22

Hence, the earnings per share is 3.22

Working Notes:

1.

Computation of common shares on January 1 to December 31 is:

Common shares on January 1 to December 31 = Common shares * (1 + Percentage of common dividend) * (Number of months from January 1 to December 31 / Total Number of months in a year)

= 680,000 * (1 + 0.04) * (12/12)

= 680,000 * 1.04 * 1

= 707,200

Hence, the common shares on January 1 to December 31 is 707,200.

2.

Computation of common shares on February 28 to December 31 is:

Common shares on February 28 to December 31 = Common shares * (1 + Percentage of common dividend) * (Number of months from February 28 to December 31 / Total Number of months in a year)

= 54,000 * (1 + 0.04) * (10/12)

= 54,000 * 1.04 * (10/12)

= 46,800

Hence, the common shares on February 28 to December 31 is 46,800.

3.

Computation of number of repurchased shares on July 1 to December 31 is:

Number of repurchased shares on July 1 to December 31 = Number of repurchased shares * (Number of months from July 1 to December 31 / Total Number of months in a year)

= 4,000 * (6/12)

= 2,000

Hence, the number of repurchased shares on July 1 to December 31 is 2,000.

4.

Computation of number of weighted shares is:

Number of weighted shares = Common shares on January 1 to December 31 + Common shares on February 28 to December 31 - Number of repurchased shares on July 1 to December 31

= 707,200 + 46,800 - 2,000

= 752,000

Hence, the number of weighted shares is 752,000.